Schedule of Cash Collections of Accounts Receivable

Pet Stop Inc., a pet wholesale supplier, was organized on May 1. Projected sales for each of the first three months of operations are as follows:

May $280,000
June 360,000
July 510,000

All sales are on account. Of sales on account, 60% are expected to be collected in the month of the sale, 35% in the first month following the sale, and the remainder in the second month following the sale.

Prepare a schedule indicating cash collections from sales for May, June, and July.

PET STOP INC.
Schedule of Collections from Sales
For the Three Months Ending July 31
May June July
May sales on account:
Collected in May
Collected in June
Collected in July
June sales on account:
Collected in June
Collected in July
July sales on account:
Collected in July
Total cash collected $ $ $
Answer #1

Prepare a schedule indicating cash collections from sales for May, June, and July.

PET STOP INC.
Schedule of Collections from Sales
For the Three Months Ending July 31
May June July
May sales on account:
Collected in May 280000*60% = 168000
Collected in June 280000*35% = 98000
Collected in July 280000*5% = 14000
June sales on account:
Collected in June 360000*60% = 216000
Collected in July 360000*35% = 126000
July sales on account:
Collected in July 510000*60% = 306000
Total cash collected $168000 $314000 $446000